# supplement

### v1

@@ 1 Main
This work considers advances that address the challenges of simulating realistic material surfaces with these quantum many-body methods, as well as using these developments to validate the accuracy and enable key insights into a chemically diverse range of outstanding surface absorption problems. This advance in simulation capabilities leverages the growing field of quantum embedding to introduce a controllable locality approximation that achieves a practical linear scaling in computational effort for system sizes considered [24]. Specifically, we extend the previously proposed ‘systematically improvable quantum embedding’ (SIE) method [25, 26], which itself builds on density matrix embedding theory and fragmentation methods of quantum chemistry [27, 28, 21, 29]. Our developments extend the approach to couple together layers of different resolutions of correlated effects at different length scales, up to the ‘gold standard’ CCSD(T) level. We demonstrate exceptional accuracy with linear scaling, crucial for performing these simulations at the required scale. We also adapt the approach to utilize graphics processing units (GPUs) to eliminate key computational bottlenecks in our workflow, including the implementation of GPU-enhanced correlated solvers. With these, we can converge unprecedented simulations at the CCSD(T) level over solid-state systems of tens of thousands of orbitals, as summarized by the workflow of Fig. 1a and the observed computational scaling of Fig. 1b, with more details of these methodological advances described in Sec. 4 and Supporting Information (SI) Sec. LABEL:sec:SI_method.
@@ 2.2 Carbonaceous molecules on various surfaces
Carbonaceous molecules are ubiquitous in nature and synthetic materials. The adsorption of these carbonaceous molecules plays a crucial role in various fields, such as pollution control, drug design and catalyst optimization [64, 65, 66, 67, 68]. As a key quantity in determining surface chemistry, adsorption energies have been computationally probed for various carbonaceous molecules and surfaces to understand their behavior [60, 57, 69, 61, 63, 62, 70, 45, 71, 43, 72, 73, 74, 75, 76]. To demonstrate the wide applicability of our method, we consider three representative systems with chemically diverse surfaces selected: carbon monoxide on the MgO(001) surface, denoted as ‘CO@MgO’, six organic molecules on coronene, a simplified model to mimic graphene, denoted as ‘Organic molecules@Coronene’ [60], and carbon monoxide [61]/carbon dioxide [63, 62] on the metal-organic framework CPO-27-Mg, denoted as ‘CO/CO2@CPO-27-Mg’. Fig. 4a-c shows the adsorption configurations and the surfaces/clusters. The adsorption energies across these systems range from $4$ to $14$ kcal/mol. This range spans from typical weak physisorption to strong chemisorption, highlighting the diversity of these distinct surface types.
@@ Acknowledgments
We extend our gratitude to Dr. Hang Li and ByteDance Research for their invaluable support. We also thank Dr. Hongzhou Ye for his insightful discussions on the project. Additionally, we appreciate Dr. Kaido Sillar and Dr. Joachim Sauer for generously sharing the CPO-27-Mg structure. J.C. is supported by the National Key R&D Program of China under Grant No. 2021YFA1400500 and the National Science Foundation of China under Grant No. 12334003.

### v2

@@ 1 Main
This work considers advances that address the challenges of simulating realistic material surfaces with these quantum many-body methods, as well as using these developments to validate the accuracy and enable key insights into a chemically diverse range of outstanding surface absorption problems. This advance in simulation capabilities leverages the growing field of quantum embedding to introduce a controllable locality approximation that achieves a practical linear scaling in computational effort for system sizes considered [24]. Specifically, we extend the previously proposed ‘systematically improvable quantum embedding’ (SIE) method [25, 26], which itself builds on density matrix embedding theory and fragmentation methods of quantum chemistry [27, 28, 21, 29, 30, 31]. Our developments extend the approach to couple together layers of different resolutions of correlated effects at different length scales, up to the ‘gold standard’ CCSD(T) level. We demonstrate exceptional accuracy with linear scaling, crucial for performing these simulations at the required scale. We also adapt the approach to utilize graphics processing units (GPUs) to eliminate key computational bottlenecks in our workflow, including the implementation of GPU-enhanced correlated solvers. With these, we can converge unprecedented simulations at the CCSD(T) level over solid-state systems of tens of thousands of orbitals, as summarized by the workflow of Fig. 1a and the observed computational scaling of Fig. 1b, with more details of these methodological advances described in Sec. 4 and Supporting Information (SI) Sec. LABEL:sec:SI_method.
@@ 2.2 Carbonaceous molecules on various surfaces
Carbonaceous molecules are ubiquitous in nature and synthetic materials. The adsorption of these carbonaceous molecules plays a crucial role in various fields, such as pollution control, drug design and catalyst optimization [66, 67, 68, 69, 70]. As a key quantity in determining surface chemistry, adsorption energies have been computationally probed for various carbonaceous molecules and surfaces to understand their behavior [62, 59, 71, 63, 65, 64, 72, 47, 73, 45, 74, 75, 76, 77, 78, 79]. To demonstrate the wide applicability of our method, we consider three representative systems with chemically diverse surfaces selected: carbon monoxide on the MgO(001) surface, denoted as ‘CO@MgO’, six organic molecules on coronene, a simplified model to mimic graphene, denoted as ‘Organic molecules@Coronene’ [62], and carbon monoxide [63]/carbon dioxide [65, 64] on the metal-organic framework CPO-27-Mg, denoted as ‘CO/CO2@CPO-27-Mg’. Fig. 4a-c shows the adsorption configurations and the surfaces/clusters. The adsorption energies across these systems range from $4$ to $14$ kcal/mol. This range spans from typical weak physisorption to strong chemisorption, highlighting the diversity of these distinct surface types.
@@ Acknowledgments
We extend our gratitude to Dr. Hang Li and ByteDance Research for their invaluable support. We thank Dr. Hong-Zhou Ye for his insightful discussions on the project. We thank Dr. Kaido Sillar and Dr. Joachim Sauer for generously sharing the CPO-27-Mg structure. Additionally, we appreciate Dr. Weiluo Ren, Weizhong Fu and Yibo Wu for their suggestions for this article. J.C. is supported by the National Key R&D Program of China under Grant No. 2021YFA1400500 and the National Science Foundation of China under Grant No. 12334003.